[Hallc_running] [New Logentry] Hall C RC update, Fri 8/31
mack at jlab.org
mack at jlab.org
Fri Aug 31 09:40:04 EDT 2018
According to Arne's guidance, Hall C shifts are cancelled for one week, until Fri Sept 7, 08:00 . Dave Meekins will watch the target today, park it over the weekend so it doesn't need continuous observation, then continue to look for fan problems next week. People who are planning to fly in the next week or so should check in with the RC at the time (Simona starts next Wed).
We basically know what we knew 24 hours ago, but people have had more coffee and more time to think. Today they are going to start to track down the leak in the cryomodule, then they will begin the time-consuming warm up to remove whatever ice has accumulated. Arne's suggested single week of cancelled shifts is the optimistic scenario where, shortly after the 3 day weekend, a simple leak is found and fixed.
We will have a Hall C RC meeting today at 4:15 pm: https://hallcweb.jlab.org/wiki/index.php/RC_daily_meetings_and_daily_runplan_fall_2018 .
After today's Hall C RC meeting, there will be no Hall C RC meetings until Tuesday Sept 4 when we will discuss ongoing work to get even ready-er for beam. The wish list so far is: 1) continue target fan tests, 2) improve big BPM/dump viewer readiness for the SHMS beam steering test, 3) expand capability to remotely change spectrometer angles, and 4) Mark J would like a missing spectrometer survey at small angles.
This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/3593186
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Hallc_running